OpenStudy (anonymous):

Which value of would make this inequality true? 12 ÷ > 5 A. 2 B. 3 C. 4 D. 6 medal and fan

OpenStudy (anonymous):

@katkipe @lala

OpenStudy (anonymous):

@M4thM1nd

OpenStudy (anonymous):

Test all the options: 12/2 is greater than 5? 12/3 is greater than 5? 12/4 is greater than 5? 12/6 is greater than 5?

OpenStudy (anonymous):

...?

OpenStudy (anonymous):

the sign \[>\] means greater than. For example \[2>1\] True

OpenStudy (anonymous):

a?

OpenStudy (secret-ninja):

12/2=6 12/3=4 12/4=3 12/6=2 So, which of those numbers is bigger than 5?

OpenStudy (secret-ninja):

Good job! :D